### v1

# Simultaneous analysis of approximate leave-one-out cross-validation and mean-field inference

Abstract.
Approximate Leave-One-Out Cross-Validation (ALO-CV) is a method that has been proposed to estimate the generalization error of a regularized estimator in the high-dimensional regime where dimension and sample size are of the same order, the so called “proportional regime”. A new analysis is developed to derive the consistency of ALO-CV for non-differentiable regularizer under Gaussian covariates and strong-convexity of the regularizer. Using a conditioning argument, the difference between the ALO-CV weights and their counterparts in mean-field inference is shown to be small. Combined with upper bounds between the mean-field inference estimate and the leave-one-out quantity, this provides a proof that ALO-CV approximates the leave-one-out quantity as well up to negligible error terms. Linear models with square loss, robust linear regression and single-index models are explicitly treated.
: Rutgers University. Email: pierre.bellec@rutgers.edu
1. Introduction to ALO-CV and mean-field inference
Consider iid observations $(x_{i},y_{i})_{i=1,...,n}$ with $x_{i}\in\mathbb{R}^{p}$ and $y_{i}$ valued in some fixed set $\mathcal{Y}$. For a test function of interest $g:\mathbb{R}\times\mathcal{Y}\to\mathbb{R}$, we are interested in the generalization error with respect to $g$, namely
$\operatorname{Err}_{g}(b)=\mathbb{E}\Bigl{[}g(b^{T}x_{new},~{}y_{new})\Bigr{]},$
for any candidate vector $b\in\mathbb{R}^{p}$, where $(x_{new},y_{new})$ is an independent copy of $(x_{i},y_{i})$. The goal of the present paper is to analyse simultaneously two statistical methods to estimate the generalization error of an estimate
$\hat{b}=\operatorname*{arg\,min}_{b\in\mathbb{R}^{p}}\sum_{i=1}^{n}L_{y_{i}}(x_{i}^{T}b)+R(b)$ (1.1)
for convex loss functions $L_{y_{i}}:\mathbb{R}\to\mathbb{R}$ and a convex penalty $R:\mathbb{R}^{p}\to\mathbb{R}$. In this case, the generalization error can be written as the conditional expectation
$\operatorname{Err}_{g}(\hat{b})=\mathbb{E}\Bigl{[}g(\hat{b}^{T}x_{new},~{}y_{new})\mid(x_{i},y_{i})_{i=1,...,n}\Bigr{]}.$ (1.2)
Above and throughout the paper, $n$ is the sample size and $p$ the dimension. The ratio $p/n$ is kept of constant order as $n,p\to+\infty$, with
$\delta\leq\frac{p}{n}\leq\delta^{-1}$ (1.3)
for some constant $\delta\in(0,1)$. This is commonly referred to as the proportional regime, of proportional asymptotics, which has been the focus of much recent research in high-dimensional statistics and optimization. We refer to [24] for an introduction to the topic. Let us survey two methods that have been proposed to estimate (1.2) in the regime (1.3).
1.1. Approximate leave-one-out cross-validation
A recent series of works [25, 34, 33, 26, 1, 35] have proposed and analyzed the Approximate Leave-One-Out Cross-Validation (ALO-CV) method, which we now review. The starting point is the leave-one-out estimate
$b^{i}=\operatorname*{arg\,min}_{b\in\mathbb{R}^{p}}\sum_{l\in[n]\setminus\{i\}}L_{y_{l}}(x_{l}^{T}b)+R(b),$ (1.4)
where the $i$-th observation is left out compared to (1.1). The leave-one-out estimate of (1.2) is then given by
$\operatorname{LOO}\coloneq\frac{1}{n}\sum_{i=1}^{n}g(x_{i}^{T}b^{i},y_{i}).$ (1.5)
Error bounds of the estimation error of (1.2) by the leave-one-out estimate (1.5) and consistency are obtained in [34, 35]. However, the leave-one-out estimate (1.5) requires to solve the minimization problem $n$ times. ALO-CV attempts to approximate (1.5) while solving the minimization problem only once on the full dataset. To do so, the ALO-CV estimate [25] is constructed by performing one Newton step, starting from $\hat{b}$, to minimize (1.4). The resulting approximation of $b^{i}$ is
$\tilde{b}_{\textsc{newton}}^{i}=\hat{b}+\Bigl{(}\sum_{k\in[n]\setminus\{i\}}x_{k}L_{y_{k}}^{\prime\prime}(x_{k}^{T}\hat{b})+\nabla^{2}R(\hat{b})\Bigr{)}^{-1}x_{i}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})$ (1.6)
which is well-defined as long as $L_{y_{k}}$ is twice differentiable with second derivative $L_{y_{k}}^{\prime\prime}$ and the Hessian $\nabla^{2}R$ exists and is positive-definite at $\hat{b}$. Replacing $b^{i}$ by $\tilde{b}_{\textsc{newton}}^{i}$ in the leave-one-out estimate (1.5), and using The Woodbury-Sherman-Morrison matrix identity to obtain
$x_{i}^{T}\Bigl{(}\sum_{k\in[n]\setminus\{i\}}x_{k}L_{y_{k}}^{\prime\prime}(x_{k}^{T}\hat{b})+\nabla^{2}R(\hat{b})\Bigr{)}^{-1}x_{i}=\frac{x_{i}^{T}\hat{A}x_{i}}{1-x_{i}^{T}\hat{A}x_{i}L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})}$
where $\hat{A}=(\sum_{i=1}^{n}x_{i}L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})x_{i}^{T}+\nabla^{2}R(\hat{b}))^{-1}$. This leads to the ALO-CV estimate of (1.2) given by
$\operatorname{ALO}\coloneq\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})\frac{x_{i}^{T}\hat{A}x_{i}}{1-x_{i}^{T}\hat{A}x_{i}L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})},~{}y_{i}\Bigr{)}.$ (1.7)
Each step above is detailed in [25, 1]. This construction has been extended to some regularizer that are not differentiable such as $R(b)=\lambda\|b\|_{1}$ [33, 32, 1, 25, §2.2].
ALO-CV is appealing because its construction is simply based on the leave-one-out estimate and the above Newton step, with no assumption on the underlying data-generating process. It is computationally efficient compared to the leave-one-out estimate (1.5). It does not suffer from the sample size bias seen when using 5- or 10-fold cross-validation to approximate (1.2) [34, 25, Figure 1]: Estimating (1.2) by 5-fold cross-validation produces approximations the generalization error of an estimator trained on $4/5$ of the data, which is not the same as the generalization error of (1.1) trained on the full dataset.
Proving the success ALO-CV for estimating (1.5) or (1.2), for instance consistency, was initially obtained for Gaussian designs under differentiability and smoothness conditions on the loss functions $L_{y_{i}}$ and the penalty $R$ [25, 34]. The differentiability and smoothness conditions were recently relaxed to allow for the elastic-net penalty under stability conditions on the supports $S^{i}=\{j\in[p]:b^{i}_{j}\neq 0\}$ of (1.4) [1, Theorem 3.2], and such stability condition is shown to hold for the square loss $L_{y_{i}}(t)=(y_{i}-t)^{2}/2$ and the elastic-net penalty under isotropic Gaussian $x_{i}$ [1, §4]. The work [1] leverages these support stability results and smoothing argument to prove consistency of ALO-CV.
This paper presents a completely different argument, without relying on smoothing arguments or support stability conditions. The main contribution of the paper is a simple argument to prove the consistency of ALO-CV for non-smooth regularizer under Gaussian design, by directly relating the ALO-CV estimate (1.7) to results obtained in mean-field asymptotics that we now describe.
1.2. Mean-field asymptotics
The last decade has seen the development of mean-field asymptotics, which aims to study the asymptotic behavior of high-dimensional estimators such as (1.1) in the proportional regime (1.3), see for instance [17, 2, 19, 29, 16, 18, 30, 23, 14, 22] and [24] for an introduction to the topic. The picture that emerges is that the behavior of the empirical distribution of $\hat{b}$ or of $(x_{i}^{T}\hat{b})_{i\in[n]}$ can be precisely described by a few deterministic scalars that are the solution of a system of nonlinear equations. The precise definition of these deterministic quantities is not relevant for the purpose of the present paper and the aforementioned papers provide numerous examples. For the purpose of the present paper, let us mention that under regularity conditions on some test function $g$ of interest, if $x_{i}\sim N(0,I_{p})$ and $y_{i}\mid x_{i}$ follows a single-index model in the sense that $y_{i}=F(x_{i}^{T}w,E_{i})$ for some deterministic $F(\cdot,\cdot)$ and $w\in\mathbb{R}^{p}$ with $\|w\|_{2}=1$ as well as some external randomness $E_{i}$ independent of $x_{i}$, these results typically provide under Gaussian design and suitable assumptions
$\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+\gamma_{*}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b}),y_{i}\Bigr{)}\to^{P}\mathbb{E}\Bigl{[}g(a_{*}U+\sigma_{*}G,y)\Bigr{]}$ (1.8)
where $(\gamma_{*},a_{*},\sigma_{*})$ are deterministic scalars found by solving the aforementioned system of equations, and in the right-hand side $(U,y)$ is equal in distribution to $(x_{i}^{T}w,y_{i})$ and $G\sim N(0,1)$ is independent of $(U,y)$. See for instance [22, Theorem 2]. Such approximation could be the starting point in deriving estimates of the generalization error (1.2), although in practice $(\gamma_{*},a_{*},\sigma_{*})$ are unknown: they need to be estimated from the data, or at the very least approximations such as (1.8) need to hold with $\gamma_{*}$ replaced by an observable quantity $\hat{\gamma}$. Estimating $(\gamma_{*},a_{*},\sigma_{*})$ with observable quantities has been the focus of [3] by the author, where observable quantities that estimate $(\gamma_{*},a_{*},\sigma_{*})$ are developed and referred to as observable adjustment. We describe in the next subsection a version of (1.8) where $\gamma_{*}$ in the left-hand side is replaced by an observable quantity $\hat{\gamma}$, which is a key intermediate ingredient for the analysis of ALO-CV of the present paper.
1.3. Mean-field inference
One family of results in mean-field asymptotics focuses on inference regarding the components $\beta_{j}^{*}$ of a ground-truth $\beta^{*}$ in regression models, or of the components $w_{j}$ of a ground-truth $w$ in single-index models. This is referred to as de-biasing, where the initial estimate $\hat{b}_{j}$ provided by the $j$-th component of (1.1) requires a bias correction after which the de-biased estimate is approximately normal and centered at $\beta_{j}^{*}$ in regression [8, 9, 14], or at a shrunk version of or $w_{j}$ in single index models [3, Section 4.1].
Another family of results useful for inference lets us construct estimate of the generalization error (1.2). The principle here is similar to de-biasing: for a given coordinate $i\in[n]$, start with the initial value $x_{i}^{T}\hat{b}$ and find a de-biasing correction that makes $x_{i}^{T}\hat{b}+\textsf{correction}$ approximately normal, see for instance [3, Theorem 4.3]. Let us state formally one version of such results that will be useful for the present paper.
We will use the notation $X\in\mathbb{R}^{n\times p}$ for the matrix with rows $x_{1},...,x_{n}$ and
$D=\operatorname{diag}(L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})_{i\in[n]})\in\mathbb{R}^{n\times n}.$ (1.9)
Furthermore, $e_{i}\in\mathbb{R}^{n}$ is the $i$-th canonical basis vector in $\mathbb{R}^{n}$, and $e_{j}\in\mathbb{R}^{p}$ the $j$-th canonical basis vector in $\mathbb{R}^{p}$. We say that the regularizer is $(\mu n,\Sigma)$-strongly convex if $b\mapsto R(b)-\mu n\|\Sigma^{1/2}b\|_{2}^{2}/2$ is convex, or equivalently
$\inf_{u\in\partial R(b)}\inf_{u^{\prime}\in\partial R(b^{\prime})}(b-b^{\prime})^{T}(u-u^{\prime})\geq n\mu\|\Sigma^{1/2}(b-b^{\prime})\|_{2}^{2},$ (1.10)
Proposition 1.1.
Consider iid $(x_{i},y_{i})_{i\in[n]}$ with $x_{i}\sim N(0,\Sigma)$, and $w\in\mathbb{R}^{p}$ with $\mathbb{E}[(w^{T}x_{i})^{2}]\in\{0,1\}$ such that $(I_{p}-\Sigma ww^{T})x_{i}$ is independent of $(x_{i}^{T}w,y_{i})$. Assume that for all value of $y\in\mathcal{Y}$, the loss $L_{y}(\cdot)$ is differentiable and 1-Lipschitz, with $L_{y}^{\prime}(\cdot)$ being 1-Lipschitz. Assume that the regularizer $R$ is $(n\mu,\Sigma)$-strongly convex in the sense of (1.10) and that $R$ is minimized at 0. Then for fixed values of $(y_{i})_{i\in[n]}$ the mapping
$(x_{ij})_{i\in[n],j\in[p]}\mapsto\hat{b}$
is almost everywhere differentiable with $\frac{\partial}{\partial x_{ij}}\hat{b}=\hat{A}(-e_{j}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})-X^{T}De_{i}\hat{b}_{j})$ for some invertible matrix $\hat{A}\in\mathbb{R}^{p\times p}$ with
$\|\Sigma^{1/2}\hat{A}^{-1}\Sigma^{1/2}\|_{op}\leq(n\mu)^{-1}.$ (1.11)
Furthermore, with the leave-one-out estimate $b^{i}$ in (1.4),
$\sum_{i=1}^{n}\mathbb{E}\Bigl{[}\Bigl{(}x_{i}^{T}\hat{b}+\operatorname{tr}[\Sigma\hat{A}]L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})-x_{i}^{T}b^{i}\Bigr{)}^{2}\Bigr{]}\leq C(\mu,\delta)$ (1.12)
for some constant depending on $\mu,\delta$ only provided that $p/n$ satisfies (1.3). If additionally $g$ satisfies $\sup_{y\in\mathcal{Y}}|g(x,y)-g(x,y^{\prime})|\leq|x-x^{\prime}|(1+|x|+|x^{\prime}|)$ then
$\mathbb{E}\Big{|}\operatorname{LOO}-\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+\operatorname{tr}[\Sigma\hat{A}]L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b}),~{}y_{i}\Bigr{)}\Big{|}\leq\frac{C(\mu,\delta)}{\sqrt{n}}.$ (1.13)
Since $x_{i}\sim N(0,I_{p})$ is independent of $b^{i}$, (1.12) shows that $x_{i}^{T}b^{i}$ plus the additive correction $\operatorname{tr}[\Sigma\hat{A}]L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})$ is approximately normal; informally
$x_{i}^{T}\hat{b}+\operatorname{tr}[\Sigma\hat{A}]L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})\approx N(0,1)\|b^{i}\|_{2}$
and $\|b^{i}\|_{2}$ in the right-hand side can be replaced by $\|\hat{b}\|_{2}$ up to a small error term by (1.14) below. The statement of Proposition 1.1 and its proof are variants of Theorem 4.3 in [3] with a minor variation on how the normal random variable is constructed. Here the standard normal random variable $x_{i}^{T}b^{i}/\|b^{i}\|_{2}$ is obtained by leaving one observation out, in [3, Theorem 4.3] it is obtained by replacing the $i$-th row of $X\in\mathbb{R}^{n\times p}$ by an independent copy. Since the results are not exactly, let us provide a short proof.
Proof of Proposition 1.1.
Let us prove the case $\Sigma=I_{p}$. The case $\Sigma\neq I_{p}$ can be deduced from the case $\Sigma=I_{p}$ by a change of variable as explained in [5, Appendix B] or [3, Appendix C.2].
The derivative formula, the existence of the matrix $\hat{A}$ and the upper bound on $\|\hat{A}\|_{op}$ are proved in [3, Proposition 3.1] (or in [5] in linear regression). Let $P=I_{p}-ww^{T}$. Applying the consequence [7, (2.6)] of the Second Order Stein formula to the Gaussian vector $Px_{i}$ conditionally on $(x_{i}^{T}w,y_{i})$ gives
$\mathbb{E}\Bigl{[}B_{i}^{2}\Bigr{]}\leq\mathbb{E}\Bigl{[}\|P(\hat{b}-b^{i})\|_{2}^{2}+\sum_{j=1}^{p}\|\frac{\partial\hat{b}}{\partial x_{ij}}\|_{2}^{2}\Bigr{]}\text{ where }B_{i}=x_{i}^{T}P\hat{b}-\sum_{j=1}^{p}e_{j}^{T}P\frac{\partial\hat{b}}{\partial x_{ij}}-x_{i}^{T}Pb^{i}.$
Using the derivative formula to bound the second term,
$\sum_{i=1}^{n}\sum_{j=1}^{p}\|\frac{\partial\hat{b}}{\partial x_{ij}}\|_{2}^{2}\leq 2\|\hat{A}\|_{F}^{2}\sum_{i=1}^{n}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})^{2}+2\|\hat{A}X^{T}D\|_{F}^{2}\|\hat{b}\|_{2}^{2}.$
For the first term, we use [1, Lemma 3 item 4][18, Appendices 2-3]:
$n\mu\|\hat{b}-b^{i}\|_{2}\leq\|x_{i}\|_{2}|L_{i}(x_{i}^{T}\hat{b})|.$ (1.14)
A short proof is as follows: write the KKT conditions $\sum_{l=1}^{n}x_{l}L_{y_{l}}^{\prime}(x_{l}^{T}\hat{b})\in-\partial R(\hat{b})$ for $\hat{b}$, the KKT conditions $\sum_{l\neq i}x_{l}L_{y_{l}}^{\prime}(x_{l}^{T}b^{i})\in-\partial R(b^{i})$ for $b^{i}$, and use the strong convexity property $(\hat{b}-b^{i})(\partial R(\hat{b})-\partial R(b^{i}))\geq n\mu\|\hat{b}-b^{i}\|_{2}^{2}$:
$n\mu\|\hat{b}-b^{i}\|_{2}^{2}$ $\leq(\hat{b}-b^{i})(\partial R(\hat{b})-\partial R(b^{i}))$
$=(\hat{b}-b^{i})^{T}x_{i}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})-\sum_{l\neq i}\Bigl{(}x_{l}^{T}\hat{b}^{T}-x_{l}^{T}b^{i}\Bigr{)}\Bigl{(}L_{y_{l}}^{\prime}(x_{l}^{T}\hat{b})-L_{y_{l}}^{\prime}(x_{l}^{T}b^{i})\Bigr{)}.$
See [4, Section 4.2] where similar arguments are used to establish smoothness properties of $\hat{b}$. Since the rightmost sum is positive by convexity of $L_{y_{l}}$, (1.14) holds by the Cauchy-Schwarz inequality. Following the idea in [4, (C.3)], writing the strong convexity property of $R$ at $\hat{b}$ and 00 for the element of $\partial R(\hat{b})$ given by the KKT conditions further gives $n\mu\|\hat{b}-0\|_{2}^{2}\leq(\hat{b}-0)(\partial R(\hat{b})-\partial R(0))=\hat{b}^{T}\sum_{i=1}^{n}x_{i}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{)}$ so that for instance
$n\mu\|\Sigma^{1/2}\hat{b}\|_{2}\leq\|X\Sigma^{-1/2}\|_{op}\sqrt{n}\max_{l\in[n]}|L_{y_{l}}(x_{l}^{T}\hat{b})|,$ (1.15)
by the same argument $n\mu\|\Sigma^{1/2}b^{i}\|_{2}\leq\|X\Sigma^{-1/2}\|_{op}\sqrt{n}\max_{l\in[n]\setminus\{i\}}|L_{y_{l}}(x_{l}^{T}b^{i})|.$ (1.16)
Let $C_{i}=x_{i}^{T}\hat{b}+\operatorname{tr}[\hat{A}]L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})-x_{i}^{T}b^{i}$ as in the left-hand side of the desired result (1.12). Using the derivative formula, with $U_{i}=x_{i}^{T}w$, we find
$C_{i}-B_{i}$ $=U_{i}w^{T}(\hat{b}-b^{i})+\operatorname{tr}[ww^{T}\hat{A}]L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})-\hat{b}^{T}P\hat{A}x_{i}L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b}).$
By the triangle inequality $\mathbb{E}[\sum_{i=1}^{n}(C_{i}+D_{i})^{2}]^{1/2}\leq\mathbb{E}[\sum_{i=1}^{n}C_{i}^{2}]^{1/2}+\mathbb{E}[\sum_{i=1}^{n}D_{i}^{2}]^{1/2}$ several times, we find that $\mathbb{E}[\sum_{i=1}^{n}C_{i}^{2}]^{1/2}$ is bounded from above by
$\mathbb{E}\Bigl{[}\sum_{i=1}^{n}B_{i}^{2}\Bigr{]}^{\frac{1}{2}}+\mathbb{E}\Bigl{[}\sum_{i=1}^{n}(U_{i}w^{T}(\hat{b}-b^{i}))^{2}\Bigr{]}^{\frac{1}{2}}+\mathbb{E}\Bigl{[}\|\hat{A}\|_{op}^{2}\sum_{i=1}^{n}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})^{2}\Bigr{]}^{\frac{1}{2}}+\mathbb{E}\Bigl{[}\|\hat{b}^{T}P\hat{A}XD\|_{F}^{2}\Bigr{]}^{\frac{1}{2}}.$
Combining these upper bounds with (1.11), (1.14), (1.15), inequalities $|L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})|\leq 1$ and $|L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})|\leq 1$ (so that $\|D\|_{op}\leq 1$) by assumption on $L_{y_{i}}$, the fact that $\mathbb{E}[\|X\|_{op}^{2}]\leq(\sqrt{n}+\sqrt{p})^{2}+1$ thanks to [15, Theorem II.13], equality $\mathbb{E}[\sum_{i=1}^{n}U_{i}^{2}\|x_{i}\|_{2}^{2}]\leq n(p-1)+3n$ since $x_{i}\sim N(0,I_{p})$ and $U_{i}=x_{i}^{T}w$, the claim (1.12) is proved. For (1.13), by assumption on the test function $g$ the left-hand side is bounded from above by $\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}|C_{i}|(1+|x_{i}^{T}b^{i}|+|C_{i}+x_{i}^{T}b^{i}|)$ so that the Cauchy-Schwarz inequality combined with (1.12) and (1.16) gives (1.13). ∎
1.4. Examples of matrices $\hat{A}$
For concreteness, let us provide the expression of $\hat{A}$ in a few examples of interest. Assuming that $L_{y_{i}}(\cdot)$ is twice-differentiable, and let $D\in\mathbb{R}^{n\times n}$ be the diagonal matrix (1.9) with $D_{ii}=L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})$.
Example 1.1.
If $R$ is twice differentiable with Hessian $\nabla^{2}R(b)$ have smaller eigenvalue bounded from below by $\nu n$ for some $\nu\geq 0$, for every $b\in\mathbb{R}^{p}$. Here, $\hat{A}=(X^{T}DX+\nabla^{2}R(\hat{b}))^{-1}.$
Example 1.2.
If $R(b)=\lambda\|b\|_{1}+n\nu\|b\|_{2}^{2}/2$ is the elastic-net for $\lambda,n\nu\geq 0$,
$\hat{A}_{\hat{S},\hat{S}}=(X_{\hat{S}}^{T}DX_{\hat{S}}+n\nu I_{\hat{S},\hat{S}})^{-1},\qquad\hat{A}_{j,k}=0\text{ if }j\notin\hat{S}\text{ or }k\notin\hat{S}$
almost surely where $\hat{S}=\{j\in[p]:\hat{b}_{j}\neq 0\}$ is the support of $\hat{b}$. Above $\hat{X}_{\hat{S}}$ is the submatrix of $X$ with columns indexed in $\hat{S}$ and $I_{\hat{S},\hat{S}}$ is the identity matrix of size $|\hat{S}|$.
Example 1.3 (group-lasso).
If $\lambda_{1},...,\lambda_{k}>0\nu\geq 0$, a partition $(G_{1},...,G_{K})$ of $[p]$ is given and $R(b)=\sum_{k=1}^{K}\lambda_{k}\|b_{G_{k}}\|_{2}+n\nu\|b\|_{2}^{2}/2$ then with $\hat{T}=\{k\in[K]:\|\hat{b}_{G_{k}}\|_{2}>0\}$ and $\hat{S}=\cup_{k\in\hat{T}}G_{K}$, almost surely
$\hat{A}_{\hat{S},\hat{S}}=\Bigl{(}X_{\hat{S}}^{T}DX_{\hat{S}}+n\nu I_{\hat{S},\hat{S}}+\sum_{k\in\hat{T}}\frac{\lambda_{k}}{\|\hat{b}_{G_{k}}\|_{2}}\Bigl{(}I_{G_{k},G_{k}}-\frac{\hat{b}_{G_{k}}\hat{b}_{G_{k}}^{T}}{\|\hat{b}_{G_{k}}\|_{2}^{2}}\Bigr{)}\Bigr{)}^{-1}$
and $\hat{A}_{j,k}=0$ if $j\notin\hat{S}$ or $k\notin\hat{S}$.
These formulae are obtained by differentiating the KKT conditions. For instance, in the case of twice-differentiable $R$, differentiating $\sum_{l=1}^{n}x_{l}L_{y_{l}}^{\prime}(x_{l}^{T}\hat{b})+\nabla R(\hat{b})=0$ with respect to $x_{ij}$ gives $e_{j}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})+(\nabla^{2}R(\hat{b})+X^{T}DX)\frac{\partial}{\partial x_{ij}}\hat{b}+X^{T}De_{i}\hat{b}_{j}=0$. The derivative formula in Proposition 1.1 thus holds with $\hat{A}$ given by Example 1.1. More care is needed in the case of Examples 1.2 and 1.3 as the penalty is non-differentiable. In such non-differentiable case, the above formulae are obtained by first showing that almost surely the active set $\hat{S}$ is locally constant; once this is established, the KKT conditions can be differentiated as in the twice-differentiable case. We refer to [9, Section 4] for such arguments applied to the lasso and group-lasso. In Proposition 1.1 the matrix $\hat{A}$ is shown to exist, however no general closed-form expression for $\hat{A}$ is available for general regularizer $R$ at this point. The above cases are the most common though, and the formulae for $\hat{A}$ in these cases are explicit and given above.
For ALO-CV, the construction with the Newton step (1.6) leading to (1.7) is only valid when the Hessian $\nabla^{2}R$ exists and is positive-definite. In this case matrix $\hat{A}$ coincides with the expression in Example 1.1. For non-differentiable penalty functions $R$ such as the L1 or elastic-net penalty, the ALO-CV construction can be extended by approximation [25, Section 2.2] leading to a well-defined matrix $\hat{A}$ in (1.7). In all cases where both
  • a closed-form formula for the matrix $\hat{A}$ in the expression $\frac{\partial}{\partial x_{ij}}\hat{b}$ in Proposition 1.1 is available, and
  • the ALO-CV construction (1.7) has been extended for some matrix $\hat{A}$,
then the expression of $\hat{A}$ in the derivative formula of Proposition 1.1 and the expression of $\hat{A}$ in the ALO-CV construction (1.7) coincide.
In existing analysis of ALO-CV for non-differentiable regularizer [1] where the penalty is the elastic-net, the major difficulty comes from controlling the difference between the support $\hat{S}$ of $\hat{b}$ defined in Example 1.2 and the support $S^{i}$ of the leave-one-out vector $b^{i}$ in (1.4). Auddy et al. [1] manages to achieve such control between the supports $S^{i}$ and $\hat{S}$ for the square loss, elastic-net penalty and isotropic Gaussian designs, and as a consequence obtains the consistency of ALO-CV in this case. This paper develops an alternative argument to derive the consistency of ALO-CV for non-differentiable regularizer, by relating ALO-CV directly to mean-field approximations such as (1.13), without trying to control the difference between the supports $\hat{S}$ and $S^{i}$ or other discrepancies between $\hat{b}$ and $b^{i}$.
1.5. Reconciling ALO-CV and mean-field inference
An observation that motivated the present study is that ALO-CV provides at first glance a different picture than the usual mean-field asymptotics results of the proportional regime. ALO-CV has been shown [25, 34, 1] to consistently estimate the leave-one-out estimate (1.5), i.e.,
$\operatorname{ALO}=\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})W_{i},~{}y_{i}\Bigr{)}\approx\operatorname{LOO}\text{ where }W_{i}\coloneq\frac{x_{i}^{T}\hat{A}x_{i}}{1-x_{i}^{T}\hat{A}x_{i}L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})}$
under suitable assumptions. On the other hand, Proposition 1.1 shows that
$\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})\operatorname{tr}[\Sigma\hat{A}],~{}y_{i}\Bigr{)}\approx\operatorname{LOO}.$ (1.17)
Typical proportional asymptotics results such as (1.8) (e.g., Theorems 1 and 2 in [22]) also suggest that the weight in front of $L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})$ should be independent of $i$ to obtain the mean-field limit of the generalization error [22, Theorem 1, eq. (2.5)].
Mean-field inference thus points to the weight $\operatorname{tr}[\Sigma\hat{A}]$ in (1.17) while ALO-CV points to the weights $W_{i}$ above. This raises the question of whether the weights $(W_{i})_{i\in[n]}$ concentrate around $\operatorname{tr}[\Sigma\hat{A}]$ or the deterministic value $\gamma_{*}$ in (1.8) (it is shown that $\operatorname{tr}[\Sigma\hat{A}]-\gamma_{*}\to^{P}0$ in [20, Lemma 13] in regression for separable loss and penalty, so that $\operatorname{tr}[\Sigma\hat{A}]$ and $\gamma_{*}$ can be used interchangeably).
Relating ALO-CV to estimates of the form (1.17) (with a single weight independent of $i$) has appeared before in [34]. In [34, equation (17)], the ALO-CV weights $W_{i}$ are shown, under suitable conditions that include twice-differentiability of both the loss and penalty, to concentrate uniformly in $i\in[n]$ around a data-driven scalar (independent of $i$) found by solving a nonlinear equation [34, (8)-(9)]). However, the argument in [34] relies on twice-differentiability of the loss and penalty and requires that the second derivatives are themselves $\alpha$-Holder continuous for some $\alpha\in(0,1]$.
When comparing the weight $\operatorname{tr}[\Sigma\hat{A}]$ in (1.17) to the weights $W_{i}$ in ALO-CV, a resembling known approximation that comes to mind is
$\operatorname{tr}[\Sigma\hat{A}]\approx\frac{\operatorname{tr}[X\hat{A}X^{T}D]}{\operatorname{tr}[D-DX\hat{A}X^{T}D]}=\frac{\sum_{i=1}^{n}x_{i}^{T}\hat{A}x_{i}L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})}{\sum_{i=1}^{n}(1-x_{i}^{T}\hat{A}x_{i}L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b}))L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b})},$ (1.18)
see [5, Section 5] in linear regression and [3] in single-index models. Without the two sums in the numerator and the denominator, the rightmost fraction would be the same as the $W_{i}$ weights from ALO-CV. The approximation (1.18) in [5, 3] is proved for nondifferentiable regularizers, so (1.18) provides a little hope towards reconciling ALO-CV with (1.17) under no smoothness assumption on the regularizer $R(\cdot)$ in (1.1). However, the proofs of (1.18) in [3, 5] do leverage the two summations in the numerator and denominator of (1.18). The gist of the present paper is a novel argument, different from the leave-one-out bounds in [34, 1] or the Stein formulae argument to prove (1.18) in [5, 3], to obtain $W_{i}\approx\operatorname{tr}[\Sigma\hat{A}]$ for non-differentiable regularizers.
1.6. Notation
[…]
2. Square loss and rotational invariance
Let us first consider the square loss $L_{y_{i}}(t)=(y_{i}-t)^{2}/2$ in (1.1), so that the minimization problem becomes
$\hat{b}=\operatorname*{arg\,min}_{b\in\mathbb{R}^{p}}\|y-Xb\|_{2}^{2}+R(b),$ (2.1)
and assume that $(X,y)$ has continuous distribution. It is known that for any fixed $X$, the mapping $y\mapsto X\hat{b}$ is 1-Lipschitz (see, e.g., [6]) so that its Jacobian
$H\coloneq\frac{\partial}{\partial y}X\hat{b}\in\mathbb{R}^{n\times n}$
exists almost everywhere. The matrix $H$ is sometimes referred to the “hat” matrix and appears in the literature on Stein’s unbiased risk estimate [28]. In this case, the matrix $\hat{A}$ in Equation 1.3 (whose existence is granted by [5, Theorem 1] for strongly-convex regularizer) is related to $H$ by $H=X\hat{A}X^{T}$.
In this case, the ALO-CV weights are given by
$W_{i}=\frac{H_{ii}}{1-H_{ii}}.$
On the other hand, [5, Section 5] proves that under a $n\mu$-strong-convexity assumption on $R$ and (1.3),
$\mathbb{E}\Bigl{[}\Big{|}\operatorname{tr}[\Sigma\hat{A}]-\frac{\operatorname{tr}[H]}{n-\operatorname{tr}[H]}\Big{|}\Bigr{]}\leq\frac{C(\mu,\delta)}{\sqrt{n}}$ (2.2)
so that $\operatorname{tr}[\Sigma\hat{A}]$ and $\operatorname{tr}[H]/(n-\operatorname{tr}[H])$ can be used interchangeably in results of the form (1.13) or (1.17). The following result shows that the approximation $W_{i}\approx\operatorname{tr}[H]/(n-\operatorname{tr}[H])$ holds uniformly in $i\in[n]$.
Proposition 2.1.
Consider the regularized least-squares estimate (2.1) where $R:\mathbb{R}^{p}\to\mathbb{R}$ is convex, and assume a linear model $y=X\beta^{*}+\varepsilon$ where $\varepsilon\sim N(0,\sigma^{2})$ is independent of $X$ and the rows $x_{i}$ of $X$ are iid $N(0,\Sigma)$. Then
$\max_{i\in[n]}|H_{ii}-\operatorname{tr}[H]/n|\leq c\sqrt{\log(n)/n}$ (2.3)
with probability at least $1-c/n$ for some absolute constant $c>0$. Furthermore, if $R$ in (2.1) is $(\mu n,I_{p})$-strongly convex then in the same event
$\Big{|}W_{i}-\frac{\operatorname{tr}[H]}{n-\operatorname{tr}[H]}\Big{|}=\Big{|}\frac{H_{ii}}{1-H_{ii}}-\frac{\operatorname{tr}[H]}{n-\operatorname{tr}[H]}\Big{|}\leq c\sqrt{\log(n)/n}\Bigl{(}1+\|X\|_{op}^{2}/(n\mu)\Bigr{)}^{2}.$ (2.4)
With (2.2) and (2.4), we see that
$\operatorname{tr}[\Sigma\hat{A}],\qquad\frac{\operatorname{tr}[H]/n}{1-\operatorname{tr}[H]/n},\qquad W_{i}=\frac{H_{ii}}{1-H_{ii}},$
can be used interchangeably in results of the form (1.17) up to a small error term of order $\sqrt{\log(n)/n}$. The proof given below relies mostly on the rotational invariance of the square loss and the Gaussian distribution. Smoothness of the regularizer is irrelevant. Inequality (2.3) only requires convexity of the regularizer, while for (2.4) we additionally assume strong-convexity as this is a simple condition that lets us control $1/(1-\operatorname{tr}[H]/n)$. Controlling $1/(1-\operatorname{tr}[H]/n)$ without strong-convexity is possible, e.g., [14], however this requires more involved arguments and different assumptions such as small-enough sparsity of $\beta^{*}$ [14, 4].
Let us explain with an example how (2.3) can be used with existing results in mean-field asymptotics to draw conclusions about ALO-CV cross-validation. By Theorems 7 and 8 in [14], for any 1-Lipschitz function $\phi:\mathbb{R}^{n}\to\mathbb{R}$, if the sparsity of $\beta^{*}$ is small enough (in the sense of the phase transition studied in [14]), the Lasso $\hat{b}$ satisfies the convergence in probability
$\phi\Bigl{(}\frac{y-X\hat{b}}{(1-\operatorname{tr}[H]/n)\sqrt{n}}\Bigr{)}-\mathbb{E}[\phi(\tau_{*}Z)]\to^{P}0$ (2.5)
where $Z\sim N(0,1)$ and $\tau_{*}^{2}$ is such that of $\sigma^{2}+\|\Sigma^{1/2}(\hat{b}-\beta^{*})\|_{2}^{2}-\tau_{*}^{2}\to^{P}0$ in the linear model $y=X\beta^{*}+\varepsilon$ where $X$ has iid $N(0,\Sigma)$ rows and $\varepsilon\sim N(0,\sigma^{2}I_{n})$ is independent of $X$. Consider a test function $g:\mathbb{R}^{2}\to\mathbb{R}$ in (1.2) of the form $g(a,y)=\varphi(a-y)$, so that the ALO-CV estimate (1.7) is
$\operatorname{ALO}=\frac{1}{n}\sum_{i=1}^{n}\varphi\Bigl{(}x_{i}^{T}\hat{b}-y_{i}+W_{i}(x_{i}^{T}\hat{b}-y_{i})\Bigr{)}=\frac{1}{n}\sum_{i=1}^{n}\varphi\Bigl{(}(1+W_{i})(x_{i}^{T}\hat{b}-y_{i})\Bigr{)}$
and $1+W_{i}=1/(1-H_{ii})$. Theorem 8 in [14] shows that $\operatorname{tr}[H]/n=\|\hat{b}\|_{0}/n$ is bounded away from 1 in the sense that $01-\operatorname{tr}[H]/n>c_{0}>$ 0 with high-probability, so that by (2.3) the scalar $H_{ii}$ is also bounded away from 1 uniformly in $i\in[n]$. Since $\varphi$ is Lipschitz, (2.3) gives
$\Big{|}\operatorname{ALO}-\frac{1}{n}\sum_{i=1}^{n}\varphi\Bigl{(}\frac{x_{i}^{T}\hat{b}-y_{i}}{1-\operatorname{tr}[H]/n}\Bigr{)}\Big{|}\lesssim\sqrt{\log(n)/n}$
with high-probability. Choosing $\phi(u)=\frac{1}{n}\sum_{i=1}^{n}\varphi(u_{i}\sqrt{n})$ in (2.5) which is 1-Lipschitz by the Cauchy-Schwarz inequality gives
$\Big{|}\operatorname{ALO}-\mathbb{E}[\varphi(\tau_{*}Z)]\Big{|}.$
Since $\sigma^{2}+\|\Sigma^{1/2}(\hat{b}-\beta^{*})\|_{2}^{2}-\tau_{*}^{2}\to^{P}0$ by [14], the quantity $\mathbb{E}[\varphi(\tau_{*}Z)]$ is the same as (1.2) up to a negligible error term converging to 0 in probability. This proves rigorously that if the sparsity of $\beta^{*}$ is small enough in the sense of the phase transition in [14], then ALO-CV consistently estimates the generalization error (1.2) for test functions of the form $g(a,y)=\varphi(a-y)$ for 1-Lipschitz functions $\varphi$ (a natural choice being $\varphi(u)=|u|$).
Proof of Proposition 2.1.
Let us realize $(X,\varepsilon)$ as $(QG,Qz)$ where $Q\in\mathbb{R}^{n\times n}$ is random orthogonal matrix in $O(n)$ distributed according to the Haar measure, and $(G,z)$ are independent $G$ having iid $N(0,\Sigma)$ rows and $z\sim N(0,\sigma^{2})$. Then $\hat{b}$ in (2.1) is independent of $Q$
$\hat{b}=\operatorname*{arg\,min}_{b\in\mathbb{R}^{p}}\|y-Xb\|_{2}^{2}+R(b)=\operatorname*{arg\,min}_{b\in\mathbb{R}^{p}}\|z-Gb\|_{2}^{2}+R(b)$ (2.6)
due to the square loss which gives $\|y-Xb\|_{2}=\|Qz-QGb\|_{2}=\|z-Gb\|_{2}$ for all $b\in\mathbb{R}^{p}$. Furthermore, with $\bar{y}=z+G\beta^{*}=Q^{T}y$ and $\bar{H}=\frac{\partial}{\partial\bar{y}}G\hat{b}$ we have by the chain rule
$H$ $=\tfrac{\partial}{\partial y}(X\hat{b})$ by definition of $H$
$=Q\tfrac{\partial}{\partial y}(G\hat{b})$ by linearity
$=Q\tfrac{\partial}{\partial\bar{y}}(G\hat{b})Q^{T}$ by the chain rule
$=Q\bar{H}Q^{T}.$
By construction, $\bar{H}$ is independent of $Q$ since this is the hat matrix in the linear model $\bar{y}=G\beta^{*}+z$ of the estimator with design matrix $G$ and response $\bar{y}$. Furthermore, $\bar{H}$ is symmetric with eigenvalues in $[0,1]$ by [7, Remark 3.3]. Since $Q^{T}e_{i}$ for $e_{i}\in\mathbb{R}^{n}$ the $i$-th canonical basis vector is uniformly distributed on the sphere, by the Hanson-Wright inequality (see for instance [27, Proposition 6.4 with $t=\log(n^{2})$]) we find since $\bar{H}$ is symmetric with eigenvalues in $[0,1]$ that
$\forall i\in[n],\qquad\mathbb{P}[|H_{ii}-\operatorname{tr}[\bar{H}]/n|\leq c\sqrt{n/\log n}+c\log(n)/n]\geq 1-c/n^{2}$
for some absolute constant $c>0$. The union bound over $i\in[n]$ proves (2.3) since $\operatorname{tr}\bar{H}=\operatorname{tr}H$. By the argument [5, between inequalities (D.7) and (D.8)], if $R$ is $(n\mu,I_{p})$-strongly convex we have $\|H\|_{op}\leq 1/((n\mu)/\|X\|_{op}^{2}+1)$, so that using $H_{ii}\leq\|H\|_{op}$ and $\operatorname{tr}[H]/n\leq\|H\|_{op}$ we get
$\max\Bigl{\{}\frac{1}{1-\operatorname{tr}[H]/n},\frac{1}{1-H_{ii}}\Bigr{\}}\leq\frac{(n\mu)/\|X\|_{op}^{2}+1}{(n\mu)/\|X\|_{op}^{2}}=1+\|X\|_{op}^{2}/(n\mu).$
Due to $|\frac{H_{ii}}{1-H_{ii}}-\frac{\operatorname{tr}H/n}{1-\operatorname{tr}[H]/n}|\leq\frac{|H_{ii}-\operatorname{tr}[H]/n|}{1-H_{ii}}+\frac{\operatorname{tr}[H]/n|H_{ii}-\operatorname{tr}[H]/n|}{(1-H_{ii})(1-\operatorname{tr}[H]/n)}$ this proves (2.4). ∎
3. Beyond rotational invariance: probabilistic results
Beyond the square loss, we cannot use rotational invariance as in the previous section. The probabilistic result that lets us control the error in the approximation $W_{i}\approx\operatorname{tr}[\hat{A}\Sigma]$ is the following.
Theorem 3.1.
Let $\mu>0$ and $\delta$ as in (1.3). Let $X$ had iid $N(0,\Sigma)$ rows with $\Sigma$ invertible. Consider matrices $P\in\mathbb{R}^{p\times p}$ and $Q\in\mathbb{R}^{n\times n}$ such that $\Sigma^{1/2}P\Sigma^{-1/2}$ and $Q$ are orthogonal projections, deterministic $\bar{U}\in\mathbb{R}^{n\times p}$ and $\bar{V}\in\mathbb{R}^{n\times p}$ and the event
$E=\Bigl{\{}X(I_{p}-P)=\bar{U},\qquad(I_{n}-Q)X=\bar{V}\Bigr{\}}.$ (3.1)
Let $H\in\mathbb{R}^{p\times p}$ with $H-n\mu I_{p}$ positive semi-definite. Let $D\in\mathbb{R}^{n\times n}$ be diagonal with entries in $[0,1]$. Let $A=(X^{T}DX+\Sigma^{1/2}H\Sigma^{1/2})^{-1}$ and define
$\operatorname{Rem}_{i}=x_{i}^{T}Ax_{i}-\operatorname{tr}[A\Sigma](1-D_{ii}x_{i}^{T}Ax_{i}).$ (3.2)
Then the constants $K,K^{\prime}$ from Proposition A.1 satisfy
$K^{\prime}\leq\frac{\delta^{-1/2}}{\mu\sqrt{n}}\max\{1,\frac{\delta^{-1/2}}{\sqrt{\mu}}\},\qquad K\leq\frac{4}{\delta\mu^{3/2}\sqrt{n}}+\frac{2}{\sqrt{\delta}\mu^{3/2}n},\qquad K+K^{\prime}\leq\frac{C(\delta,\mu)}{\sqrt{n}},$
and we have for some absolute constant $C>0$, all $t\geq C$ and any fixed $i\in[n]$
$\mathbb{P}\Bigl{(}\Big{|}\operatorname{Rem}_{i}-\mathbb{E}[\operatorname{Rem}_{i}\mid E]\Big{|}>tK+tK^{\prime}\Bigr{)}\leq 2\exp(-t^{2}/2)+C\exp(-t^{2/3}/C).$ (3.3)
With $d_{P}=\dim\ker P=\operatorname{rank}(I_{p}-P)$ and $d_{Q}=\dim\ker Q=\operatorname{rank}(I_{n}-Q)$, the conditional expectations satisfy
$\mathbb{E}\Bigl{[}\sum_{i=1}^{n}\mathbb{E}[\operatorname{Rem}_{i}\mid E]^{2}\Bigr{]}\leq c\Bigl{(}\frac{d_{P}+d_{Q}}{\delta}+\frac{1}{n^{2}\mu^{2}}\Bigr{)}\frac{1}{\delta\mu^{2}}+c\Bigl{(}\frac{d_{P}^{2}}{\mu^{2}n}+\frac{d_{Q}}{\delta^{2}\mu^{2}}\Bigr{)}\Bigl{(}1+\frac{1}{\delta\mu}\Bigr{)}$ (3.4)
for some absolute constant $c>0$. The right-hand side is smaller than a constant $C(\delta,\mu,d_{P},d_{Q})$ depending only on $\delta,\mu,d_{P},d_{Q}$.
The proof is given in Appendix A In other words, $\operatorname{Rem}_{i}$ concentrates exponentially fast (although with $t^{3/2}$ inside the exponential) around its conditional expectation $\mathbb{E}[\operatorname{Rem}_{i}\mid E]$. Using a union bound over $i\in[n]$ taking $t^{2/3}=C\log(n^{2})$ in (3.3), we get with probability $1-C/n$ that
$\max_{i\in[n]}|\operatorname{Rem}_{i}-\mathbb{E}[\operatorname{Rem}_{i}\mid E]|\leq c(K+K^{\prime})\log(n)^{3/2},\qquad$
for some absolute constant $c>0$ and $(K,K^{\prime})$ are of order $n^{-1/2}$. The conditional expectations are themselves controlled since $\sum_{i=1}^{n}\mathbb{E}[\operatorname{Rem}_{i}\mid E]^{2}$ is bounded in expectation by a constant when $\mu,\delta,d_{P},d_{Q}$ are kept constant as in the applications below. For instance, if $S=\sum_{i=1}^{n}\mathbb{E}[\operatorname{Rem}_{i}\mid E]^{2}$ then
|$\Big{|}\{i\in[n]:\mathbb{E}[\operatorname{Rem}_{i}\mid E]^{2}<\frac{S}{\sqrt{n}}\}\Big{|}\geq n-\sqrt{n}$
so that all but $\sqrt{n}$ conditional expectations $(\mathbb{E}[\operatorname{Rem}_{i}\mid E])_{i\in[n]}$ are of order $S/\sqrt{n}=O_{P}(n^{-1/2})$. Different than the exponential concentration argument obtained above, a different strategy would be to use [7, (2.6)] to control $\sum_{i=1}^{n}\mathbb{E}[\operatorname{Rem}_{i}^{2}\mid E]$ directly. This leads to a better dependence on $n$ in the upper bound, but the exponential concentration around $\mathbb{E}[\operatorname{Rem}_{i}\mid E]$ is lost. We discuss this approach in Appendix B.
To obtain $W_{i}\approx\operatorname{tr}[\Sigma\hat{A}]$ from the above, the last ingredient we will need is that $1-D_{ii}x_{i}^{T}Ax_{i}$ in the definition of $\operatorname{Rem}_{i}$ (3.2) is bounded away from 0. This follows, with the notation and assumptions in Theorem 3.1 from
$x_{i}^{T}Ax_{i}\leq x_{i}^{T}(x_{i}D_{i}x_{i}^{T}+n\mu\Sigma)^{-1}x_{i}=\frac{x_{i}^{T}\Sigma^{-1}x_{i}}{D_{ii}x_{i}^{T}\Sigma^{-1}x_{i}+n\mu}$ (3.5)
by comparing positive definite matrices inside the inverse, and solving the linear system with right-hand side $x_{i}$ using
$(x_{i}D_{i}x_{i}^{T}+n\mu\Sigma)\frac{\Sigma^{-1}x_{i}}{D_{i}x_{i}^{T}\Sigma^{-1}x_{i}+n\mu}=x_{i}.$
Multiplying by $-D_{ii}$ and adding 1, inequality (3.5) is equivalent to
$1-D_{ii}x_{i}^{T}Ax_{i}\geq 1-\frac{D_{ii}x_{i}^{T}\Sigma^{-1}x_{i}}{D_{ii}x_{i}^{T}\Sigma^{-1}x_{i}+n\mu}=\frac{n\mu}{D_{ii}x_{i}^{T}\Sigma^{-1}x_{i}+n\mu},$
so that $1-D_{ii}x_{i}^{T}Ax_{i}$ is bounded away from 0 in the sense
$0<(1-D_{ii}x_{i}^{T}Ax_{i})^{-1}\leq 1+D_{ii}x_{i}^{T}\Sigma^{-1}x_{i}/(n\mu).$ (3.6)
Restriction to a support or a subset
In order to apply the above result to Examples 1.2 and 1.3, it is useful to state the following corollary, where
$A=\lim_{t\to+\infty}(X^{T}DX+\Sigma^{1/2}H\Sigma^{1/2}+tW)^{-1}$ (3.7)
for some orthogonal projection $W\in\mathbb{R}^{n\times n}$. If $S\subset[p]$ is a fixed subset of indices and $W=\sum_{j\in S}e_{j}e_{j}^{T}$, the above limit is given by
$A_{S,S}=(X_{S}^{T}DX_{S}+(\Sigma^{1/2}H\Sigma^{1/2})_{S,S})^{-1},\qquad A_{jk}=0\text{ if }j\notin S\text{ or }k\notin S.$ (3.8)
Corollary 3.2.
Let the assumptions of Theorem 3.1 be fulfilled. Let $W$ be a deterministic orthogonal projection. Replace the definition of $A$ by (3.7) in (3.1). Then the conclusions (3.3), (3.4) and (3.6) still hold.
With $A_{t}=(X^{T}DX+\Sigma^{1/2}H\Sigma^{1/2}+tW)^{-1}$ for $t\geq 0$, we have the monotone limits $\operatorname{tr}[\Sigma^{1/2}A_{t}\Sigma^{1/2}]\downarrow\operatorname{tr}[\Sigma^{1/2}A\Sigma^{1/2}]$ and $x_{i}^{T}A_{t}x_{i}\downarrow\operatorname{tr}[A]$ as $t\uparrow+\infty$. Applying Theorem 3.1 to $A_{t}$ and taking $t\to+\infty$ using the dominated convergence theorem gives Corollary 3.2.
What is the catch?
The notation and setting used so far in Section 3 is seemingly detached from the notation in Examples 1.2 and 1.3 or Proposition 1.1. In the result Theorem 3.1, the matrices $D$ and $H$ are deterministic, and the support $S$ in (3.8) is also deterministic for Corollary 3.2 to hold with the deterministic projection $W=\sum_{j\in S}e_{j}e_{j}^{T}$.
In the setting discussed in the introduction, the corresponding quantities are random. With Example 1.1, we are interested in
$\hat{A}=(X^{T}DX+\nabla^{2}R(\hat{b}))^{-1}$
which is random because $\nabla^{2}R(\hat{b})$ is random through $\hat{b}$, and $D=\operatorname{diag}(\{L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b}),i\in[n]\})$ is also random through $L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b}),i\in[n]$. In Examples 1.2 and 1.3, the support $\hat{S}=\{j\in[p]:\hat{b}_{j}\neq 0\}$ is random as well, and it is unclear at this point why Theorem 3.1 and Corollary 3.2 with a deterministic $S\subset[p]$ in (3.8) would be useful at all. This will be resolved by considering particular events of the form (3.1), such that in this event $D,\hat{b},\nabla^{2}R(\hat{b})$, and in the case of Examples 1.2 and 1.3 the support $\hat{S}$, are all fully determined (i.e., they are conditionally conditionally on an event of the form (3.1)). Such events have been studied in [12, 13] and are described in the following sections.
4. Conditioning example I: Robust linear regression
Before we move to generalized linear models and single-index models in Section 5, let us apply the main result in Theorem 3.1 to robust linear regression, where a linear model
$y_{i}=x_{i}^{T}\beta^{*}+\varepsilon_{i}$ (4.1)
is assumed, with $x_{i}\sim N(0,I_{p})$ and $\varepsilon_{i}$ is independent of $X$. Consider throughout this section a differentiable robust loss function $\rho:\mathbb{R}\to\mathbb{R}$ such as the Huber loss, so that the minimization problem is
$\hat{b}=\operatorname*{arg\,min}_{b\in\mathbb{R}^{p}}\sum_{i=1}^{n}\rho(y_{i}-x_{i}^{T}\hat{b})+R(b)$ (4.2)
The key intuition behind the argument of Proposition 2.1 is that in the square loss case, if $Q\Lambda P^{T}$ is the SVD of the design matrix $[X\mid\varepsilon]\in\mathbb{R}^{n\times(p+1)}$, then $\hat{b}$ and $\operatorname{tr}[H]$ are independent of the left singular vectors $Q$. We can then argue conditionally on $(\Lambda,P)$ with respect to the randomness of the randomness of the independent matrix $Q$.
This argument fails as soon as the loss $L_{y_{i}}(\cdot)$ is not the square loss: the rotational invariance argued in (2.6) is lost. Instead, we argue by conditioning on the following event. Let us argue conditionally on the noise $\varepsilon$ (which is independent of $X$), so that $\varepsilon$ is fixed. The only remaining randomness comes from the design matrix $X$ that we assume has iid $N(0,\Sigma)$ rows. Consider deterministic $\bar{h}\in\mathbb{R}^{p},\bar{u}\in\mathbb{R}^{n},\bar{v}\in\mathbb{R}^{n},\bar{w}\in\mathbb{R}^{p}$ with
$\bar{w}\in\partial R(\bar{h}+\beta^{*}),\qquad\bar{v}_{i}=\rho^{\prime}(\varepsilon_{i}-\bar{u}_{i}).$ (4.3)
Then consider the event
$X\bar{h}=\bar{u},\quad X^{T}\bar{v}=\bar{w}.$ (4.4)
This event is studied in [12, 13] in order to apply the Convex Gaussian Minmax Theorem of [30] sequentially. The first key observation made in [12, 13] is that in the event (4.4), the vector $\bar{h}+\beta^{*}$ is solution to the optimization problem (4.2) because the KKT conditions
$X^{T}\bar{v}=\sum_{i=1}^{n}x_{i}\rho^{\prime}\Bigl{(}y_{i}-x_{i}^{T}(\bar{h}+\beta^{*})\Bigr{)}\in\partial R(\bar{h}+\beta^{*})$
are satisfied. Hence in this event $\hat{b}=\bar{h}+\beta^{*}$, $X(\hat{b}-\beta^{*})=\bar{u}$ and $\bar{w}$ is an element of $\partial R(\hat{b})$ that satisfies the KKT conditions for $\hat{b}$. The second key observation made in [12, 13] is that the event (4.4) is made of linear constraints on the Gaussian matrix $X$. By properties of the multivariate normal distribution, conditionally on (4.4), the entries of $X$ are again jointly normal.
The event (4.4) lets us essentially condition on $\hat{b},X(\hat{b}-\beta^{*})$ and $X^{T}\rho^{\prime}(y-X\hat{b})$, while maintaining that the entries of $X$ are jointly normal (although with different mean and covariance than the original random matrix).
Corollary 4.1.
Let $\mu,\delta>0$ be constants. Assume the proportional regime (1.3). Consider the robust regression setting (4.1)-(4.1) with $\varepsilon$ independent of $X$ and loss $\rho$ differentiable with $\rho^{\prime}$ 1-Lipschitz. Assume that $X$ has iid $N(0,\Sigma)$ rows with $\Sigma$ invertible. Consider one of the regularizer in Examples 1.1, 1.2 and 1.3 and assume $\nu\|\Sigma\|_{op}\geq\mu$ so that the strong-convexity (1.10) holds. For the explicit matrix $\hat{A}$ given in Examples 1.1, 1.2 and 1.3 we have
$\frac{1}{n}\sum_{i=1}^{n}\Big{|}W_{i}-\operatorname{tr}[\hat{A}\Sigma]\Big{|}^{2}=\frac{1}{n}\sum_{i=1}^{n}\Big{|}\frac{x_{i}^{T}\hat{A}x_{i}}{1-D_{ii}x_{i}^{T}\hat{A}x_{i}}-\operatorname{tr}[\hat{A}\Sigma]\Big{|}^{2}\leq\frac{\operatorname{polylog}_{\mu,\delta}(n)}{n}.$ (4.5)
with probability approaching one, where $\operatorname{polylog}_{\mu,\delta}$ is a polynomial in $\log(n)$ with coefficients depending on $\mu,\delta$ only. If additionally the regularizer $R(\cdot)$ is minimized at 0, the loss $\rho$ is 1-Lipschitz and the test function $g$ satisfies $\sup_{y\in\mathcal{Y}}|g(x,y)-g(x,y^{\prime})|\leq|x-x^{\prime}|(1+|x|+|x^{\prime}|)$ then
$\Big{|}\operatorname{LOO}-\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+\frac{x_{i}^{T}\hat{A}x_{i}}{1-D_{ii}x_{i}^{T}\hat{A}x_{i}}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b}),~{}y_{i}\Bigr{)}\Big{|}\leq\frac{\operatorname{polylog}_{\mu,\delta}(n)}{\sqrt{n}}.$ (4.6)
with probability converging to 1 as $n,p\to+\infty$. Above, $\operatorname{LOO}$ is defined in (1.5).
Together with the bound $|\operatorname{LOO}-\operatorname{Err}(\hat{b})|=O_{P}(n^{-1/2})$ from [35] for the quantities (1.5), (1.2), the above establishes the consistency of the ALO-CV estimates (1.7) for estimating the generalization error (1.2).
Proof.
Since $\varepsilon$ is independent of $X$, we argue conditionally on $\varepsilon$ and consider $\varepsilon$ fixed in what follows. Consider deterministic $\bar{h}\in\mathbb{R}^{p},\bar{u}\in\mathbb{R}^{n},\bar{v}\in\mathbb{R}^{n},\bar{w}\in\mathbb{R}^{p}$ such that (4.3) holds and the event (4.4). In this event, $\hat{b},D$ and $\hat{S}$ are deterministic, Choose $P,Q$ in Theorem 3.1 be projections of rank $p-1$ and $n-1$ such that (3.1) is the same as (4.4). In the case of Examples 1.2 and 1.3, apply Corollary 3.2 with $W=\sum_{j\in\hat{S}}e_{j}e_{j}^{T}$, where $\hat{S}$ is deterministic thanks to conditioning. The bound follow from (3.3), (3.4) combined with Markov’s inequality, and (3.6) combined with $\mathbb{P}(x_{i}^{T}\Sigma^{-1}x_{i}>4p)\leq\exp(-p)$ [21, Lemma 1] and the union bound over $i\in[n]$. Once the result is proved conditionally on (3.1), it remains to integrate over $(\bar{h},\bar{u},\bar{v},\bar{w},\varepsilon)$ with respect to the probability distribution of $\hat{b}-\beta^{*},X(\hat{b}-\beta^{*}),\rho(y-X\hat{b}),X^{T}\rho^{\prime}(y-X\hat{b}),\varepsilon)$.
Once (4.5) is proved, to show (4.6) we apply (1.13) and use Markov’s inequality to obtain that the integrand in the left-hand side of (1.13) is smaller than $n^{-1/4}$ with probability approaching one. By the assumed properties of $g$ we have thanks to $|L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})|\leq 1$ and the Cauchy-Schwarz inequality
$\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+\frac{x_{i}^{T}\hat{A}x_{i}}{1-D_{ii}x_{i}^{T}\hat{A}x_{i}}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b}),~{}y_{i}\Bigr{)}-\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+\operatorname{tr}[\hat{A}\Sigma]L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b}),~{}y_{i}\Bigr{)}$
(4.5)$\Bigl{(}\frac{1}{n}\sum_{i=1}^{n}1+\Bigl{(}x_{i}^{T}\hat{b}+W_{i}L_{y_{i}}(x_{i}^{T}\hat{b})\Bigr{)}^{2}+\Bigl{(}x_{i}^{T}\hat{b}+\operatorname{tr}[\Sigma\hat{A}]L_{y_{i}}(x_{i}^{T}\hat{b})\Bigr{)}^{2}\Bigr{)}^{1/2}.$
By $\operatorname{tr}[\Sigma\hat{A}]\leq\frac{p}{n\mu}$, (4.5) to bound $W_{i}$, $|L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b})|\leq 1$ and (1.15) to bound $\frac{1}{n}\|X\hat{b}\|_{2}^{2}$ we obtain (4.6). ∎
5. Conditioning example II: Single-index models
In this section, consider a single index model for $y_{i}\mid x_{i}$, of the form
$y_{i}=F(x_{i}^{T}w,\varepsilon_{i})$ (5.1)
where $\varepsilon_{i}$ is some external randomness independent of $X$, and $w\in\mathbb{R}^{p}$ is an unknown deterministic index, normalized with $\mathbb{E}[(x_{i}^{T}w)^{2}]=1$ by convention (as the amplitude of $w$ can be otherwise absorbed into $F$). In binary classification, we may for instance take $\varepsilon_{i}\sim$Unif$[0,1]$ and $y_{i}=I\{U_{i}\leq\phi(x_{i}^{T}W)\}$ where $\phi:\mathbb{R}\to[0,1]$ is the sigmoid (logistic regression) or the Gaussian CDF (probit regression).
In robust linear regression with (4.1)-(4.2), there is only direction $\bar{h}\in\mathbb{R}^{p}$ that we need to condition upon on the right of $X$ in (4.4), because in this event (and for fixed $\varepsilon$), $(y_{i},L_{y_{i}},x_{i}^{T}(\hat{b}-\beta^{*}),y_{i},\hat{b}-\beta^{*},\hat{b})$ are all deterministic. In single index model, the nonlinearity requires us to condition on another direction on the right of $X$.
By conditioning, we may consider the external randomness $\varepsilon_{i}$ fixed and argue only with respect to the randomness of $X$. Consider, for deterministic $\bar{b},\bar{u},\bar{u}^{\prime}\in\mathbb{R}^{n},\bar{v}\in\mathbb{R}^{n},\bar{h}\in\mathbb{R}^{p},\bar{y}\in\mathcal{Y}^{n}$ such that
$\bar{y}_{i}=F(\bar{u}^{\prime}_{i},\varepsilon_{i}),\qquad\bar{v}_{i}=L_{\bar{y}_{i}}^{\prime}(\bar{u}_{i}),\qquad\bar{h}\in\partial R(\bar{b})$
and the event
$E=\{X\bar{b}=\bar{u},\quad Xw=\bar{u}^{\prime},\quad X^{T}\bar{v}=\bar{h}\}.$ (5.2)
In this event, we have by construction that $y_{i}=\bar{y}_{i}$ and that $\bar{b}$ satisfies the KKT conditions of the problem (1.1), hence $\hat{b}=\bar{v}$ and $x_{i}^{T}\hat{b}=\bar{u}_{i}$. Conditionally in this event, the responses $y_{i}$, the loss $L_{y_{i}}(\cdot)$, the estimator $\hat{b}$ and its predicted values $x_{i}^{T}\hat{b}$, the matrix $D=\operatorname{diag}(\{L_{y_{i}}^{\prime\prime}(x_{i}^{T}\hat{b}),i\in[n]\})$ and the support $\hat{S}=\{j\in[p]:\hat{b}_{j}\neq 0\}$ are all fully determined and deterministic.
Corollary 5.1.
Let $\mu,\delta>0$ be constants. Assume the proportional regime (1.3). Consider the single index model (5.1) with $\varepsilon$ independent of $X$, and the estimator $\hat{b}$ in (1.1). Assume that $X$ has iid $N(0,\Sigma)$ rows with $\Sigma$ invertible. Assume that for all $y\in\mathcal{Y}$, the function $L_{y}(\cdot)$ is differentiable with $L_{y}^{\prime}(\cdot)$ 1-Lipschitz. Consider one of the regularizer in Examples 1.1, 1.2 and 1.3 and assume $\nu\|\Sigma\|_{op}\geq\mu$ so that the strong-convexity (1.10) holds. For the explicit matrix $\hat{A}$ given in Examples 1.1, 1.2 and 1.3 we have
$\frac{1}{n}\sum_{i=1}^{n}\Big{|}W_{i}-\operatorname{tr}[\hat{A}\Sigma]\Big{|}^{2}=\frac{1}{n}\sum_{i=1}^{n}\Big{|}\frac{x_{i}^{T}\hat{A}x_{i}}{1-D_{ii}x_{i}^{T}\hat{A}x_{i}}-\operatorname{tr}[\hat{A}\Sigma]\Big{|}^{2}\leq\frac{\operatorname{polylog}_{\mu,\delta}(n)}{n}.$ (5.3)
with probability approaching one. If additionally the regularizer $R(\cdot)$ is minimized at 0, the loss $y\mapsto L_{y}(\cdot)$ is 1-Lipschitz for all $y\in\mathcal{Y}$ and the test function $g$ satisfies $\sup_{y\in\mathcal{Y}}|g(x,y)-g(x,y^{\prime})|\leq|x-x^{\prime}|(1+|x|+|x^{\prime}|)$ then
$\Big{|}\operatorname{LOO}-\frac{1}{n}\sum_{i=1}^{n}g\Bigl{(}x_{i}^{T}\hat{b}+\frac{x_{i}^{T}\hat{A}x_{i}}{1-D_{ii}x_{i}^{T}\hat{A}x_{i}}L_{y_{i}}^{\prime}(x_{i}^{T}\hat{b}),~{}y_{i}\Bigr{)}\Big{|}\leq\frac{\operatorname{polylog}_{\mu,\delta}(n)}{\sqrt{n}}.$ (5.4)
with probability converging to 1 as $n,p\to+\infty$. Above, $\operatorname{LOO}$ is defined in (1.5).
The proof is exactly the same as in Corollary 4.1 with the following difference: Here we choose choose $P,Q$ in Theorem 3.1 be projections of rank $p-2$ and $n-1$ such that the event (3.1) is the same as (5.2).
Appendix A Proof of Theorem 3.1
First we need the following proposition
Proposition A.1.
Let $i\in[n]$ be fixed. Let $D\in\mathbb{R}^{n\times n}$ be a deterministic diagonal matrix with entries in $[0,1]$ and let $H\in\mathbb{R}^{p\times p}$ be a symmetric positive definite matrix. Define the functions $f,F:\mathbb{R}^{n\times p}\to\mathbb{R}$ as
$f(X)=D_{ii}e_{i}^{T}X(X^{T}DX+H)^{-1}X^{T}e_{i},\qquad F(X)=\operatorname{tr}[(X^{T}DX+H)^{-1}].$
Then $f$ is $4\|H^{-1}\|_{op}^{1/2}$-Lipschitz and $F$ is $2\sqrt{p}\|H^{-1}\|_{op}^{3/2}$-Lipschitz, both with respect to the Frobenius norm. Furthermore, for all $X$ we have $|F(X)|\leq p\|H^{-1}\|_{op}$ and $0\leq f(X)\leq 1$ hence the function
$X\mapsto F(X)(1-f(X))$ (A.1)
is $K$-Lipschitz for $K=\|f\|_{lip}p\|H\|^{-1}\|_{op}+\|F\|_{lip}\leq 4p\|H^{-1}\|_{op}^{3/2}+2\sqrt{p}\|H^{-1}\|_{op}^{3/2}$. Let $K^{\prime}\coloneq\|H^{-1}\|_{op}\sqrt{p}\max\{1,\|H^{-1}\|_{op}^{1/2}\sqrt{p}\}$. The function $h:X\mapsto\frac{f(X)}{K^{\prime}D_{ii}}=\frac{x_{i}^{T}(X^{T}DX+H)^{-1}x_{i}}{K^{\prime}}$ satisfies
$\frac{d}{dt}h(X+t\dot{X})\Big{|}_{t=0}=\frac{1}{K^{\prime}}\Bigl{(}2e_{i}^{T}\dot{X}AX^{T}e_{i}-2e_{i}^{T}XA(X^{T}D\dot{X}+\dot{X}^{T}DX)AX^{T}e_{i}\Bigr{)}$ (A.2)
which is bounded as
$\frac{1}{\|\dot{X}\|_{op}}\Big{|}\frac{d}{dt}h(X+t\dot{X})\big{|}_{t=0}\Big{|}$ $\leq\frac{2\|x_{i}\|_{2}\|H^{-1}\|_{op}+2\|x_{i}\|_{2}^{2}\|H^{-1}\|_{op}^{3/2}}{K^{\prime}}\leq 1+\frac{3\|x_{i}\|_{2}^{2}}{p}.$
Proof.
Let $A=(X^{T}DX+H)^{-1}$. Then the bounds
$\|A\|_{op}\leq\|H^{-1}\|_{op},~{}~{}\|D^{1/2}XAX^{T}D^{1/2}\|_{op}\leq 1,~{}~{}\|D^{1/2}XA\|_{op}\leq\|H^{-1}\|_{op}^{1/2}$ (A.3)
hold due to the definition of $A$. For $F$, by differentiating, for any direction $\dot{X}\in\mathbb{R}^{n\times p}$
$\frac{d}{dt}F(X+t\dot{X})\Big{|}_{t=0}=\operatorname{tr}\Bigl{[}-A(\dot{X}^{T}DX+X^{T}D\dot{X})A\Bigr{]}\leq 2\|D^{1/2}\dot{X}\|_{F}\|D^{1/2}XA^{2}\|_{F}$
and $\|D^{1/2}XA^{2}\|_{F}\leq\sqrt{p}\|D^{1/2}XA\|_{op}\|A\|_{op}\leq\sqrt{p}\|H^{-1}\|_{op}^{3/2}$. For $f$, again by differentiating,
$\frac{d}{dt}f(X+t\dot{X})\Big{|}_{t=0}=2D_{ii}e_{i}^{T}\dot{X}AX^{T}e_{i}-D_{ii}e_{i}^{T}XA(X^{T}D\dot{X}+\dot{X}^{T}DX)AX^{T}e_{i}.$
The first term is bounded from above by
$2\|D^{1/2}\dot{X}\|_{op}\|AX^{T}D^{1/2}\|_{op}\leq 2\|D^{1/2}\dot{X}\|_{op}\|H^{-1}\|_{op}^{1/2}.$
The second term is bounded from above by
$2\|D^{1/2}XAX^{T}D^{1/2}\|_{op}\|D^{1/2}\dot{X}\|_{op}\|AX^{T}D^{1/2}\|_{op}\leq 2\|D^{1/2}\dot{X}\|_{op}\|H^{-1}\|_{op}^{1/2}.$
The final claim (A.1) follows from the chain rule. ∎
We are now ready to prove Theorem 3.1.
Proof of Theorem 3.1.
Let us first treat the case $\Sigma=I_{p}$. The case $\Sigma\neq I_{p}$ will be handled later by a change of variable argument. The matrix $QXP$ is independent of $E$, and conditionally on $E$ the matrix $QXP$ has jointly normal entries with mean 0. An explicit representation is
$X=QX+\bar{V}=QXP+Q\bar{U}+\bar{V},$ (A.4)
so that conditionally on $E$, all the randomness comes from the first term $QXP$.
By Proposition A.1, conditionally on $E$ we have that $\operatorname{tr}[A\Sigma](1-D_{ii}x_{i}^{T}Ax_{i})$ is a $K$-Lipschitz function of $QXP$. Since $QXP$ is the image by a linear transformation of operator norm 1 of a standard normal vector, the concentration of Lipschitz functions of standard normal vector [11, Theorem 5.6] gives
$\mathbb{P}\Bigl{(}\Big{|}\operatorname{tr}[A\Sigma](1-D_{ii}x_{i}^{T}Ax_{i})-\mathbb{E}\Bigl{[}\operatorname{tr}[A\Sigma](1-D_{ii}x_{i}^{T}Ax_{i})\mid E\Bigr{]}\Big{|}>tK\Bigr{)}\leq 2\exp(-t^{2}/2).$
For the other term in $\operatorname{Rem}_{i}$, namely $x_{i}^{T}Ax_{i}$, we are not above at this point to get as good concentration as for the first term. Denote $h(X)=\frac{x_{i}^{T}Ax_{i}}{K^{\prime}}$ for $K^{\prime}$ as in Proposition A.1. By [10, Theorem 1.7.1], for any nondecreasing convex function $\phi$ we have
$\mathbb{P}(h(X)-\mathbb{E}[h(X)\mid E]>t\mid E)$
$\leq\mathbb{P}(\phi(h(X)-\mathbb{E}[h(X)\mid E])>\phi(t)\mid E)$ (monotonicity)
$\leq\phi(t)^{-1}\mathbb{E}[\phi(h(X)-\mathbb{E}[h(X)\mid E])\mid E]$ (Markov’s ineq.)
$\leq\phi(t)^{-1}\mathbb{E}[\phi(\tfrac{\pi}{2}\tfrac{d}{dt}h(X+t\dot{X})\big{|}_{t=0})\mid E]$ (Theorem 1.7.1 in [10])
where, conditionally on $E$, $\dot{X}$ is a independent copy of $X-\mathbb{E}[X\mid E]$; that is, from (A.4) $\dot{X}=Q\tilde{X}P$ where $\tilde{X}$ has iid $N(0,1)$ entries independent of everything else. The expression $\tfrac{d}{dt}h(X+t\dot{X})\big{|}_{t=0}$ is given by (A.2). By independence and normality of $\dot{X}$, for some $Z\sim N(0,1)$ we have $\tfrac{d}{dt}h(X+t\dot{X})\big{|}_{t=0}=ZN$ where $N$ is the L2 norm of the corresponding gradient at $X$, which is bounded from above by $1+3\|x_{i}\|_{2}^{2}/p$ thanks to Proposition A.1. We have established that for any nondecreasing convex $\phi:\mathbb{R}\to\mathbb{R}$,
$\mathbb{P}(h(X)-\mathbb{E}[h(X)\mid E]>t\mid E)\leq\phi(t)^{-1}\mathbb{E}[\phi(\tfrac{\pi}{2}|Z|(1+3\|x_{i}\|_{2}^{2}/p))\mid E]$
with $Z\sim N(0,1)$. For simplicity, let us separate $|Z|$ from $\|x_{i}\|_{2}^{2}$ in the right-hand side using Young’s inequality
$|Z|(1+3\|x_{i}\|_{2}^{2}/p)\leq|Z|+3(|Z|^{r}/r+(\|x_{i}\|_{2}^{2}/p)^{q}/q$
where $r>1,q>1$ and $11/r+1/q=$ 1 and $r=2q$, which gives $(1/2+1)/q=1$ so $q=3/2$ and $r=3$. By convexity of $\phi$ and an average of 3, we have
$\mathbb{E}[\phi(\tfrac{\pi}{2}|Z|(1+3\|x_{i}\|_{2}^{2}/p))\mid E]\leq c\mathbb{E}\Bigl{[}\phi(c|Z|)+\phi(c|Z|^{3})+\phi(c(\|x_{i}\|_{2}^{2}/p)^{3/2})\mid E\Bigr{]}$
for some absolute constant $c>0$. Now take the unconditional expectation. We need to find a convex, nondecreasing $\phi$ such that the right-hand side is bounded by a numerical constant. Choose $\phi(u)=\exp(\frac{3}{8}\max(1,\operatorname{sign}(u)|u/c|^{2/3})))$. For the second term, we are left with
$\mathbb{E}\exp(\tfrac{3}{8}\max(1,|Z|^{2}))\leq\exp(\tfrac{3}{8})\mathbb{E}\exp(\tfrac{3}{8}|Z|^{2})=\exp(\tfrac{3}{8})2$
thanks to $\mathbb{E}\exp(\tfrac{3}{8}|Z|^{2})=2$ by the explicit formula for the moment generating function of a chi-square random variable. The first term $\phi(c|Z|)$ is bounded similarly. For the third term, we have thanks to Jensen’s inequality for the average over $[p]$
$\mathbb{E}\exp(\tfrac{3}{8}\max(1,\|x_{i}\|_{2}^{2}/p))\leq e^{\tfrac{3}{8}}\mathbb{E}\exp(\tfrac{3}{8}\|x_{i}\|_{2}^{2}/p)\leq e^{\tfrac{3}{8}}\sum_{j=1}^{p}\mathbb{E}\exp(\tfrac{3}{8}x_{ij}^{2})=e^{\tfrac{3}{8}}2.$
This completes the proof that for any $t\geq 1$, for some absolute constant $C>0$,
$\mathbb{P}(h(X)-\mathbb{E}[h(X)\mid E]>t)\leq\phi(t)^{-1}C=C\exp(-(t/c)^{2/3}\tfrac{3}{8})$
and the proof of (3.3) is complete.
One version of Stein’s formula says that if $(g_{k})_{k\in[K]}$ and $g$ are jointly normal random variables with mean zero, then for any Lipschitz function $f:\mathbb{R}^{K}\to\mathbb{R}$ we have
$\mathbb{E}[(g-\mathbb{E}[g])f(g_{1},\ldots,g_{K})]=\sum_{k=1}^{K}\mathbb{E}\Bigl{[}(g_{k}-\mathbb{E}[g_{k}](g-\mathbb{E}[g])\Bigr{]}\mathbb{E}\Bigl{[}\frac{\partial}{\partial g_{k}}f(g_{1},...,g_{k})\Bigr{]}.$
With $\tilde{\mathbb{E}}$ denoting the conditional expectation given $E$, we have with $g=(e_{i}^{T}QXPe_{j})$
$\tilde{\mathbb{E}}\Bigl{[}(e_{i}^{T}QXPe_{j})e_{j}^{T}Ax_{i}\Bigr{]}=\sum_{l=1}^{n}\sum_{k=1}^{p}\tilde{\mathbb{E}}\Bigl{[}(e_{i}^{T}QXPe_{j})(x_{lk}-\tilde{\mathbb{E}}[x_{lk}])\Bigr{]}\tilde{\mathbb{E}}\Bigl{[}\frac{\partial}{\partial x_{lk}}e_{j}^{T}Ax_{i}\Bigr{]}.$
For the covariance term on the right,
$\tilde{\mathbb{E}}[(e_{i}^{T}QXPe_{j})(x_{lk}-\tilde{\mathbb{E}}[x_{lk}])]=e_{i}^{T}Qe_{l}e_{j}^{T}Pe_{k}=Q_{il}P_{jk}.$
For the term involving the derivative on the right,
$\frac{\partial}{\partial x_{lk}}e_{j}^{T}Ax_{i}=e_{j}^{T}Ae_{k}I\{i=l\}-e_{j}^{T}A\Bigl{(}e_{k}e_{l}^{T}DX+X^{T}De_{l}e_{k}^{T}\Bigr{)}Ax_{i}.$
Thus after summing over $j\in[p]$ we obtain
$\tilde{\mathbb{E}}\Bigl{[}e_{i}^{T}QXPAx_{i}\Bigr{]}=\tilde{\mathbb{E}}\Bigl{[}\operatorname{tr}[AP]e_{i}^{T}Qe_{i}-\operatorname{tr}[AP]e_{i}^{T}QDXAx_{i}-x_{i}^{T}APAX^{T}DQe_{i}\Bigr{]}.$
This is the first step to show that the expectation $\tilde{\mathbb{E}}[\operatorname{Rem}_{i}]$ is small, although there are some extra terms due to the presence of the rightmost term above, and the matrices $Q$ and $P$. We now show that these extra terms are negligible.
$\tilde{\mathbb{E}}[\operatorname{Rem}_{i}]$ $=\tilde{\mathbb{E}}\Bigl{[}e_{i}^{T}(X-QXP)Ax_{i}+e_{i}^{T}QXPAx_{i}\Bigr{]}-\tilde{\mathbb{E}}\Bigl{[}\operatorname{tr}[A](1-D_{ii}x_{i}^{T}Ax_{i})\Bigr{]}$
$=\tilde{\mathbb{E}}\Bigl{[}e_{i}^{T}(X-QXP)Ax_{i}\Bigr{]}+\tilde{\mathbb{E}}\Bigl{[}x_{i}^{T}APAX^{T}DQe_{i}\Bigr{]}$
$\quad\quad+\tilde{\mathbb{E}}\Bigl{[}\operatorname{tr}[AP]e_{i}^{T}Qe_{i}-\operatorname{tr}[A]\Bigr{]}+\tilde{\mathbb{E}}\Bigl{[}e_{i}^{T}\Bigl{(}\operatorname{tr}[A]-Q\operatorname{tr}[AP]\Bigr{)}DXAx_{i}\Bigr{]}.$
:-$\coloneq\tilde{\mathbb{E}}[A_{i}+B_{i}+C_{i}+E_{i}].:$
We now bound the four terms on the right-hand side using linear algebra. Since the sum of squares of diagonal elements of a matrix is bounded by its squared Frobenius norm, we have with $S=\sum_{i=1}^{n}A_{i}^{2}+B_{i}^{2}+C_{i}^{2}+E_{i}^{2}$ that
$S$ $\leq\|(X-QXP)AX^{T}\|_{F}^{2}+\|XAPAX^{T}DQ\|_{F}^{2}$
$\qquad+\|\operatorname{tr}[AP]Q-\operatorname{tr}[A]I_{n}\|_{F}^{2}+\|(\operatorname{tr}[A]I_{n}-\operatorname{tr}[AP]Q)DXAX^{T}\|_{F}^{2}$
$\leq\|(X-QXP)AX^{T}\|_{F}^{2}+\|XAPAX^{T}D\|_{F}^{2}+\|\operatorname{tr}[AP]Q-\operatorname{tr}[A]I_{n}\|_{F}^{2}(1+\|DXAX^{T}\|_{op}^{2}).$
thanks to $\|MM^{\prime}\|_{F}\leq\|M\|_{F}\|M\|_{op}$ that we will use repeatedly. Using $\|DXA^{1/2}\|_{op}\leq 1$ from (A.3) where possible,
$S$ $\leq\|X-QXP\|_{F}^{2}\|XA\|_{op}^{2}+\|XA\|_{op}^{2}\|A\|_{op}+\|\operatorname{tr}[AP]Q-\operatorname{tr}[A]I_{n}\|_{F}^{2}(1+\|A^{1/2}X^{T}\|_{op}^{2}).$ (A.5)
Since $X-QXP=(I_{n}-Q)X+QX(I_{p}-P)$, the rank of $X-QXP$ is at most $d_{P}+d_{Q}$ hence
$\|X-QXP\|_{F}^{2}$ $\leq(d_{P}+d_{Q})\|X-QXP\|_{op}^{2}$
$\leq(d_{P}+d_{Q})(2\|X\|_{op}^{2}+2\|QXP\|_{op}^{2})\leq 4(d_{P}+d_{Q})\|X\|_{op}^{2}.$
For the last term in the upper bound on $S$,
$\operatorname{tr}[A]I_{n}-Q\operatorname{tr}[AP]$ $=\operatorname{tr}[A](I_{n}-Q)+Q(\operatorname{tr}[A]-\operatorname{tr}[AP]),$
$\|\operatorname{tr}[A]I_{n}-Q\operatorname{tr}[AP]\|_{F}$ $\leq\|\operatorname{tr}[A](I_{n}-Q)\|_{F}+\|Q(\operatorname{tr}[A]-\operatorname{tr}[AP])\|_{F}.$
For the second term, simply use $\|Q\|_{F}\leq\sqrt{n}$ so that $\|Q(\operatorname{tr}[A]-\operatorname{tr}[AP])\|_{F}=\sqrt{n}\operatorname{tr}[A(I_{p}-P)]\leq\sqrt{n}\|A\|_{op}d_{P}$. The first term on the other hand is $\operatorname{tr}[A]\sqrt{d_{Q}}$. We have thus established the following upper bound on $S$:
$S\leq\Bigl{(}4(d_{P}+d_{Q})\|X\|_{op}^{2}+\|A\|_{op}^{2}\Bigr{)}\|XA\|_{op}^{2}+\Bigl{(}2n\|A\|_{op}^{2}d_{P}^{2}+2\operatorname{tr}[A]^{2}d_{Q}\Bigr{)}(1+\|A^{1/2}X^{T}\|_{op}^{2}).$
We use $\|A\|_{op}\leq\|H^{-1}\|_{op}\leq(n\mu)^{-1}$ and $\mathbb{E}[\|X\|_{op}^{4}]\leq c(n^{2}+p^{2})$ for some numerical constant $c$ if $X$ has iid $N(0,I_{p})$ rows (by integrating the tail bounds in [15, Theorem 2.13] or [31, Corollary 7.3.3] for instance). Combining the pieces completes the proof in the case $\Sigma=I_{p}$.
The case $\Sigma\neq I_{p}$ is obtained by considering $\tilde{X}=X\Sigma^{-1/2}$, $\tilde{Q}=Q$, $\tilde{P}=\Sigma^{1/2}P\Sigma^{-1/2}$ and $\tilde{A}=\Sigma^{1/2}A\Sigma^{1/2}$, and applying the isotropic result to $\tilde{X},\tilde{Q},\tilde{P},\tilde{A}$.
Appendix B Alternative argument to control $\sum_{i=1}^{n}\mathbb{E}[\operatorname{Rem}_{i}^{2}]$
With the notation of the proof in the previous appendix, let us mention an alternative argument to control from above $\sum_{i=1}^{n}\tilde{\mathbb{E}}[\operatorname{Rem}_{i}^{2}]$ directly, without establishing a concentration of the form (3.3). Assume $\Sigma=I_{p}$ and the setting of the proof of Theorem 3.1 in the previous appendix. The inequality given after (2.7) in [7] states that if $z\sim N(m,\Lambda)$ in $\mathbb{R}^{q}$ and $f:\mathbb{R}^{q}\to\mathbb{R}^{q}$ is locally Lipschitz then
$\mathbb{E}[((z-m)^{T}f(z)-\operatorname{tr}[\Lambda\nabla f(z)])^{2}]\leq\mathbb{E}[\|\Lambda^{1/2}f(z)\|_{2}^{2}+\|\Lambda^{1/2}\nabla f(z)\Lambda^{1/2}\|_{F}^{2}].$
Here, take $q=np$, $z$ to be the vectorization of $X=QXP+Q\bar{U}+\bar{V}$, and the function $f$ being the vectorization of $f^{i}(X)=e_{i}x_{i}^{T}A$. Then $\Lambda$ above is the Kronecker product of $Q$ and $P$ in $\mathbb{R}^{np\times np}$, which has operator norm at most one so that it can be omitted in the above upper bound. This gives $(z-m)^{T}f(z)=e_{i}^{T}QXPAx_{i}=\operatorname{tr}[f^{i}(X)^{T}QXP]$ as well as
$\operatorname{tr}[\Lambda\nabla f(z)]=\operatorname{tr}[AP]e_{i}^{T}Qe_{i}+\operatorname{tr}[AP]e_{i}^{T}QDXAx_{i}+x_{i}^{T}APAX^{T}DQe_{i},$
and with the definition
$\operatorname{Rem}_{i}^{\prime}\coloneq e_{i}^{T}QXPAx_{i}-\operatorname{tr}[AP]e_{i}^{T}Qe_{i}+\operatorname{tr}[AP]e_{i}^{T}QDXAx_{i}+x_{i}^{T}APAX^{T}DQe_{i},$
we get
$\tilde{\mathbb{E}}\Bigl{[}\Bigl{(}\operatorname{Rem}_{i}^{\prime}\Bigr{)}^{2}\Bigr{]}\leq\tilde{\mathbb{E}}\Bigl{[}\|f^{i}(X)\|_{F}^{2}+\sum_{j=1}^{p}\sum_{l=1}^{n}\|\frac{\partial}{\partial x_{lj}}f^{i}(X)\|_{F}^{2}\Bigr{]}.$
Summing over $i\in[n]$, the right-hand side and using the definition of $f^{i}$,
$\sum_{i=1}^{n}\tilde{\mathbb{E}}\Bigl{[}(\operatorname{Rem}_{i}^{\prime})^{2}\Bigr{]}\leq\tilde{\mathbb{E}}\Bigl{[}\|AX\|_{F}^{2}+\sum_{j=1}^{p}\sum_{l=1}^{n}\|\frac{\partial}{\partial x_{lj}}(AX^{T})\|_{F}^{2}\Bigr{]}.$
We have $\frac{\partial}{\partial x_{lj}}(AX^{T})=Ae_{j}e_{l}^{T}-A(e_{j}e_{l}^{T}DX+X^{T}De_{l}e_{j}^{T})AX^{T}$ by direct differentiation, hence
$\frac{1}{3}\sum_{j=1}^{p}\sum_{l=1}^{n}\|\frac{\partial(AX^{T})}{\partial x_{lj}}\|_{F}^{2}$ $\leq\sum_{j=1}^{p}\sum_{l=1}^{n}\|Ae_{j}\|_{2}^{2}+\|Ae_{j}\|_{2}^{2}\|e_{l}^{T}DXAX^{T}\|_{2}^{2}+\|AX^{T}De_{l}\|_{2}^{2}\|e_{j}^{T}AX^{T}\|_{2}^{2}$
$=n\|A\|_{F}^{2}+\|A\|_{F}^{2}\|DXAX^{T}\|_{F}^{2}+\|AX^{T}D\|_{F}^{2}\|AX^{T}\|_{F}^{2}$
$\leq n\|A\|_{F}^{2}+\|A\|_{F}^{2}\|A^{1/2}X^{T}\|_{F}^{2}+\|A^{1/2}\|_{F}^{2}\|AX^{T}\|_{F}^{2}$
$\leq n\|A\|_{F}^{2}+\|A\|_{F}^{2}\|A\|_{op}\|X^{T}\|_{F}^{2}+\operatorname{tr}[A]\|A\|_{op}^{2}\|X^{T}\|_{F}^{2}$
thanks to $\|D^{1/2}XA^{1/2}\|_{op}\leq 1$ for the last inequality. We have $\|A\|_{op}\leq(n\mu)^{-1}$ so that $\|A\|_{F}^{2}\leq p/(n\mu)^{2}$ and $\operatorname{tr}[A]\leq p/(n\mu)$ and $\mathbb{E}[\|X\|_{F}^{2}]=np$, so that the right-hand side is bounded from above in unconditional expectation by $\delta^{-1}/\mu^{2}+2\delta^{-2}/\mu^{3}$. Since $\sum_{i=1}^{n}\mathbb{E}[(\operatorname{Rem}_{i}-\operatorname{Rem}_{i}^{\prime})^{2}]$ has already been controlled by $C(\delta,\mu)$ in (A.5), this proves $\mathbb{E}[\sum_{i=1}^{n}\operatorname{Rem}_{i}^{2}]\leq C(\delta,\mu)$. Together with (3.6), this provides an alternative technique which shows that the expectation of the left-hand side of (4.5) is smaller than $C(\delta,\mu)/n$, without log factors.